# MixCache: Mixture-of-Cache for Video Diffusion Transformer Acceleration

### v1

@@
Abstract
Leveraging the Transformer architecture and the diffusion process, video DiT models have emerged as a dominant approach for high-quality video generation. However, their multi-step iterative denoising process incurs high computational cost and inference latency. Caching, a widely adopted optimization method in DiT models, leverages the redundancy in the diffusion process to skip computations in different granularities (e.g., step, cfg, block). Nevertheless, existing caching methods are limited to single-granularity strategies, struggling to balance generation quality and inference speed in a flexible manner. In this work, we propose MixCache, a training-free caching-based framework for efficient video DiT inference. It first distinguishes the interference and boundary between different caching strategies, and then introduces a context-aware cache triggering strategy to determine when caching should be enabled, along with an adaptive hybrid cache decision strategy for dynamically selecting the optimal caching granularity. Extensive experiments on diverse models demonstrate that, MixCache can significantly accelerate video generation (e.g., 1.94 speedup on Wan 14B, 1.97 speedup on HunyuanVideo) while delivering both superior generation quality and inference efficiency compared to baseline methods.
1 Introduction
Diffusion Transformer (DiT) (Peebles and Xie 2023) has revolutionized video generation by integrating the scalability of the Transformer architecture (Vaswani et al. 2017) with the power of the diffusion process (Ho, Jain, and Abbeel 2020; Rombach et al. 2022), enabling unprecedented quality in video generation. Cutting-edge video DiT models emerge including SD3.0 (Esser et al. 2024), Sora (Brooks et al. 2024), CogVideoX (Yang et al. 2024b) and Wan (Wang et al. 2025). These models have facilitated the development of many meaningful applications, such as text-to-video generation (Khachatryan et al. 2023; Wu et al. 2023), video editing (Wang et al. 2023; Jiang et al. 2025) and video continuation (Yang et al. 2024b).
@@ 1 Introduction
The research community has made significant efforts to speed up video DiT models. Among them, caching-based acceleration has emerged as one of the most widely adopted approaches. Caching leverages output similarity across diffusion timesteps by storing and reusing intermediate features, thereby reducing redundant computation and improving efficiency. According to the caching granularity from coarse to fine, existing caching methods can be divided into step level (Liu et al. 2024; Kahatapitiya et al. 2024; Ma et al. 2025), cfg level (Lv et al. 2024), block level (Ma, Fang, and Wang 2024; Wimbauer et al. 2024; Li et al. 2023; Selvaraju et al. 2024; Zhao et al. 2025; Ma et al. 2024; Shen et al. 2024), and token level (Lou et al. 2024; Zou et al. 2024b, a).
@@
2 Related Works
In the following, we introduce important works that are related to our proposed method.
Step level caching. Recent DiT advancements introduce diverse step level caching strategies. TeaCache (Liu et al. 2024) leverages input-output correlations by estimating timestep embeddings to implement a dynamic caching. AdaCache (Kahatapitiya et al. 2024) dynamically evaluates step-wise differences to determine skip lengths. AB Cache (Yu et al. 2025) extends prior methods by reusing combined outputs from the previous steps, rather than a single-step result. NIRVANA (Agarwal et al. 2024) diverges from these intra-generation approaches by enabling cross-generation cache reuse to skip computation of preceding sampleing timesteps.
CFG level caching. Recent studies leverage cfg similarities to accelerate diffusion inference. FasterCache (Lv et al. 2024) exploits the similarity between conditional and unconditional outputs at the same timestep to construct a cfg level cache mechanism. DiTFastAttn (Yuan et al. 2024) identifies consistent attention patterns in specific attention heads of conditional and unconditional inference, implementing cfg level attention sharing to bypass redundant computation.
Block level caching. For the DiT architecture, recent works propose block level caching with varying strategies. FORA (Selvaraju et al. 2024), PAB (Zhao et al. 2025), and BlockDance (Zhang et al. 2025) employ static interval-based block skipping (e.g., MLP/attention blocks). -DiT (Chen et al. 2024) adopts a phase-specific caching strategy, storing back blocks during early stages and front blocks in later stages. L2C (Ma et al. 2024) and MD-DiT (Shen et al. 2024) implement dynamic runtime caching through learnable layer selection and gradient-free search mechanisms, respectively.
Other optimizations. Recent works (Lou et al. 2024; Zou et al. 2024b, a; Liu et al. 2025) have explored token redundancy (such as background areas) to reduce computational cost. However, the dynamic nature and dataset-dependent characteristics of input tokens necessitate extensive code modifications to transformer blocks for implementation, resulting in poor compatibility. Moreover, the acceleration efficiency of token level caching is fundamentally limited by its fine-grained caching granularity and the overhead of token importance scoring. In response to the performance bottleneck caused by the multi-step iterative characteristic of the diffusion process, some works also explore efficient solvers, including DDIM (Song, Meng, and Ermon 2021), DPM-Solver (Lu et al. 2022a) and DPM-Solver++ (Lu et al. 2022b). Additionally, distillation-based methods (Salimans and Ho 2022; Luhman and Luhman 2021; Meng et al. 2023) have been developed to compress the number of diffusion sampling timesteps.
3 Methodology
3.1 Preliminary
@@
where is constant related to , and is the total sampling timesteps. A noise estimation network plays a critical role in approximating the noise distribution in the diffusion process. Specifically, it aims to minimize the discrepancy between the predicted noise term and the actual noise . In most current works, the noise estimation network adopts the DiT architecture, where the predicted noise function can be further reformulated as:
@@
Classifier-Free Guidance (CFG) has proven to be a powerful technique for improving the fidelity of generated videos in diffusion models. During the sampling process, CFG generates two distinct predictions: the conditional output conditioned on the input context , and the unconditional output derived from the empty/negative prompt . The final denoised output is obtained by:
@@ 3.2 Analysis and Motivation
Potential of three-level cache integration.
We implement different granularity caching methods at each diffusion timestep, and adopt the LPIPS (Zhang et al. 2018) metric to calculate the end-to-end similarity with the original video. The smaller the LPIPS value, the higher the similarity. As shown in Figure 3, the optimal cache granularity varies dynamically across timesteps, suggesting that adaptive cache selection could enhance generation quality.
@@
3.3 Context-aware Cache Triggering
As the early diffusion stage is responsible for overall framework sketching, it is highly sensitive to interference at this time, as described in SRDiffusion (Cheng et al. 2025). This phenomenon can also be validated in Figure 2, where in the initial diffusion process, the three-level redundancy is relatively low. Therefore, we perform full computation in the initial diffusion stage, and calculate between the output of the current timestep and that of the previous timestep , and we call this stage as the warm up stage. When is smaller than the predefined threshold , established in the offline profiling process (detailed later), the warm up phase ends and enters the cache enabled phase.
@@ 3.3 Context-aware Cache Triggering
3.4 Adaptive Hybrid Cache Decision
@@
After identifying cache enabled timesteps, it is crucial to determine the specific caching granularity for a certain timestep within the three-level caching. The limitation of prior works (Liu et al. 2024, 2025; Ma et al. 2025) is their exclusive focus on similarity of different cache methods while neglecting their differential impact on accuracy. To assess the accuracy impact of different caching methods, we generate Gaussian distribution with predefined statistical parameters (mean , standard deviation ) and calculate the distance between perturbed and original outputs. It should be noted that and should align with the real parameters derived from the three-level caching. Therefore, we firstly profile prompts and examine the and of the distance tensor between the real output and the cached output. As shown in Figure 6, the value of the distance tensor is inconsistent across different prompts, and also varies at different timesteps. However, its values are fixed within a certain range. Therefore, we set the value as the average of value across all timesteps. In addition, for the value, except for the initial timesteps, it remains stable at a fix value. Therefore, we set the value as the average of the value of the last 40 timesteps. We exclude the first 10 timesteps from analysis since these initial timesteps correspond to the warm up phase without applying caching. Consequently, their outputs lack statistical representativeness for evaluating the proposed methodology.
@@
After determining the and value, we examine the relative L1 distance between the perturbed output and the original output. We examine for all timesteps and plot the results in Figure 5(a). The smaller the distance, the smaller the impact of the interference on the actual results. This value can indicate the accuracy impact of performing a specific level cache at a specific timestep. From Figure 5(a), it can be seen that the impact value of step level and cfg level remain relatively stable after the warm up phase. However, the block level value exhibits time-dependent characteristics, where its initial value is low, and then increases in the later diffusion stage. In addition, the compare across three levels reveals that, the cfg level interference exerts a substantially greater influence, with its value exceeding both step level and block level interference by an order of magnitude. We present the visualized results and similarity metrics in Figure 5(b). It can be seen that the video quality generated using cfg level interference is poor, indicating that this interference has a significant impact on the final results, which is consistent with the conclusion in Figure 5(a). Therefore, the values presented in Figure 5(a) are employed as quantitative evaluation metrics for assessing accuracy impact, with the impact of step level and cfg level remaining constant (, ), while the block level impact exhibit as a time-dependent and index-dependent function ().
@@
In addition, to avoid getting stuck in the same caching granularity within a cache interval, we introduce penalty strategy. If a certain cache method is used in this timestep, its value in the next timestep will be multiplied by a penalty coefficient, which we set to 5 in our experiments.
3.5 Unify it together: MixCache Framework
@@
According to Figure 2, the redundancy level of different models varies. Therefore, given a specific model, MixCache will firstly undergo an offline profiling process. During the offline profiling process, MixCache profiles 100 prompts to execute the diffusion process and determines the hyper-parameters , and . In addition, the offline profiling also determines the and of random Gaussian distribution interference, and applies the corresponding interference to determine the impact value of different granularities.
After the offline profiling process, it enters the runtime inference phase. For each prompt, MixCache first applies the context-aware cache triggering strategy to determine which timestep to enable caching. For the cache enabled timesteps, MixCache uses the adaptive hybrid cache decision strategy to determine which of the three-level caching methods to use. The pseudo code of the overall execution process is presented in Appendix A.1.
4 Experiments
4.1 Settings
Base Models and Baselines. We evaluate MixCache on Wan2.1 14B (Wang et al. 2025), HunyuanVideo (Kong et al. 2024) and CogVideoX 5B (Yang et al. 2024b). For baseline methods, we choose Teacache (Liu et al. 2024), FasterCache (Lv et al. 2024), BlockDance (Zhang et al. 2025) and PAB (Zhao et al. 2025), all of which are specifically designed to accelerate DiT models through caching. Among them, Teacache adopts step level cache, FasterCache combines cfg level and block level cache, and BlockDance and PAB employ block level cache. More implementation details about baseline methods are discussed in Appendix A.2.
Evaluation Metrics. To evaluate the quality of video generation, we employ VBench (Huang et al. 2024), a widely-adopted comprehensive benchmarking suite for evaluating video generation. Based on VBench standard prompt set, we use Qwen2.5-14B-Instruct (Yang et al. 2024a) to extend all prompts to enhance the video quality, and generate 5 videos with different seeds for each prompt. In addition, we report LPIPS (Zhang et al. 2018), SSIM (Wang et al. 2004) and PSNR for quality comparison. For efficiency evaluation, we quantify the average inference latency per prompt as the primary performance metric.
Experiment Settings. For the main results, we generated 4720 videos for each set of results based on VBench. These were processed in data parallel across a public cloud instance with 64 NVIDIA A800 (80GB) GPUs, and the whole experiment takes nearly a month to complete. In the ablation study, we randomly sample 200 prompts from VBench to conduct experiments, each using 1 seed for generation.
Implementation Details. For block level caching, more block candidates will generate larger memory overhead, and even cause OOM error; Fewer block indexes will lose cache flexibility. Therefore, we pre-defined block index as 10, 20, and 30 as candidate block cache objects. More details about the offline profiling are presented in Appendix A.3.
4.2 Main Results
Quantitative Comparison. Table 1 presents a quantitative comparison of MixCache with baselines in terms of efficiency and visual quality. The results highlight that, MixCache consistently demonstrates robust acceleration efficiency and maintains visually compelling quality across diverse base models, baselines, and resolutions. Full VBench scores are shown in the Appendix.
@@ 4.2 Main Results
Visual Comparison. Figure 7 compares the videos generated by MixCache against those by the original model and baselines. The results demonstrate that MixCache can effectively preserve the original semantics and fine details. More visual results are presented in Appendix A.4.
4.3 Ablation Study
In the ablation study, we investigate the effects of scaling strategy and three-level caching respectively in terms of the visual quality and efficiency. The quantitative results are shown in Table 2 and the corresponding visualization results are presented in Figure 8. Among them, “hybrid” refers to using the adaptive hybrid cache decision strategy to determine the cache method for each timestep, while “step/cfg/block only” refers to using only a specific caching method. “N=4” represents a fixed cache interval of 4, and “” represents enabling the scaling strategy that prioritizes efficiency. It can be seen that MixCache outperforms all ablation experiments in both quality and efficiency metrics, indicating that combining dynamic scaling and three-level caching can achieve better generation quality and inference efficiency. Among them, “cfg only+” can achieve generation quality comparable to MixCache, as the cfg level redundancy is very high (reflected in Figure 1 where its relative L1 distance is small). However, due to its limited computation skip proportion, it is far inferior to MixCache in terms of efficiency.

### v2

# Adaptive Hybrid Caching for Efficient Text-to-Video Diffusion Model Acceleration

@@
Abstract.
Efficient video generation models are increasingly vital for multimedia synthetic content generation. Leveraging the Transformer architecture and the diffusion process, video DiT models have emerged as a dominant approach for high-quality video generation. However, their multi-step iterative denoising process incurs high computational cost and inference latency, which limits their practical deployment in large-scale and interactive multimedia applications. Caching, a widely adopted optimization method in DiT models, leverages the redundancy in the diffusion process to skip computations in different granularities (e.g., step, cfg, block). Nevertheless, existing caching methods are limited to single-granularity strategies, struggling to balance generation quality and inference speed in a flexible manner. In this work, we propose MixCache, a training-free caching-based framework for efficient video DiT inference. MixCache first distinguishes the interference and boundary between different caching strategies, and then introduces a context-aware cache triggering strategy to determine when caching should be enabled, along with an adaptive hybrid cache decision strategy for dynamically selecting the optimal caching granularity. Extensive experiments on diverse models demonstrate that MixCache can significantly accelerate video generation (e.g., 1.94× speedup on Wan 14B, 1.97× speedup on HunyuanVideo) while delivering both superior generation quality and inference efficiency compared to baseline methods.
1. Introduction
Diffusion Transformer (DiT) (Peebles and Xie, 2023) has revolutionized video generation by integrating the scalability of the Transformer architecture (Vaswani et al., 2017) with the power of the diffusion process (Ho et al., 2020; Rombach et al., 2022), enabling unprecedented quality in synthetic video creation. Recent advances in video DiT models, such as SD3.0 (Esser et al., 2024), Sora (Brooks et al., 2024), CogVideoX (Yang et al., 2024b), and Wan (Wang et al., 2025), have greatly expanded the capabilities of multimedia systems, supporting a wide range of applications including text-to-video generation (Khachatryan et al., 2023; Wu et al., 2023), video editing (Wang et al., 2023; Jiang et al., 2025), and video continuation (Yang et al., 2024b). Such advancements not only improve the fidelity and variety of generated multimedia content but also open new possibilities for creative expression in video-based applications.
@@ 1. Introduction
The research community has made significant efforts to speed up video DiT models. Among them, caching-based acceleration has emerged as one of the most widely adopted approaches. Caching can reduce redundant computation and improve efficiency by storing and reusing intermediate features, leveraging output similarity across diffusion timesteps. According to the caching granularity from coarse to fine, existing caching methods can be divided into step level (Liu et al., 2024; Kahatapitiya et al., 2024; Ma et al., 2025), cfg level (Lv et al., 2024), block level (Ma et al., 2024b; Wimbauer et al., 2024; Li et al., 2023; Selvaraju et al., 2024; Zhao et al., 2025; Ma et al., 2024a; Shen et al., 2024), and token level (Lou et al., 2024; Zou et al., 2024b, a).
@@
2. Related Works
In the following, we introduce works that are highly related to our method.
Step level caching. Recent DiT advancements introduce diverse step level caching strategies. TeaCache (Liu et al., 2024) leverages input-output correlations by estimating timestep embeddings to implement a dynamic caching. AdaCache (Kahatapitiya et al., 2024) dynamically evaluates step-wise differences to determine skip lengths. AB Cache (Yu et al., 2025) extends prior methods by reusing combined outputs from the previous steps, rather than a single-step result. NIRVANA (Agarwal et al., 2024) diverges from these intra-generation approaches by enabling cross-generation cache reuse to skip computation of preceding sampling timesteps.
CFG level caching. Recent studies leverage cfg similarities to accelerate diffusion inference. FasterCache (Lv et al., 2024) exploits the similarity between conditional and unconditional outputs at the same timestep to construct a cfg level cache mechanism. DiTFastAttn (Yuan et al., 2024) identifies consistent attention patterns in specific attention heads of conditional and unconditional inference, implementing cfg level attention sharing to bypass redundant computation.
Block level caching. For the DiT architecture, recent works propose block level caching with varying strategies. FORA (Selvaraju et al., 2024), PAB (Zhao et al., 2025), and BlockDance (Zhang et al., 2025) employ static interval-based block skipping (e.g., MLP/attention blocks). -DiT (Chen et al., 2024) adopts a phase-specific caching strategy, storing back blocks during early stages and front blocks in later stages. L2C (Ma et al., 2024a) and MD-DiT (Shen et al., 2024) implement dynamic runtime caching through learnable layer selection and gradient-free search mechanisms, respectively.
Other optimizations. Recent works (Lou et al., 2024; Zou et al., 2024b, a; Liu et al., 2025) have explored token redundancy (such as background areas) to reduce computational cost. However, the dynamic nature and dataset-dependent characteristics of input tokens necessitate extensive code modifications to transformer blocks for implementation, resulting in poor compatibility. Moreover, the acceleration efficiency of token level caching is fundamentally limited by its fine-grained caching granularity and the overhead of token importance scoring. In response to the performance bottleneck caused by the multi-step iterative characteristic of the diffusion process, some works also explore efficient solvers, including DDIM (Song et al., 2021), DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b). Additionally, distillation-based methods (Salimans and Ho, 2022; Luhman and Luhman, 2021; Meng et al., 2023) have been developed to compress the number of diffusion sampling timesteps.
3. Methodology
3.1. Preliminary
@@
where is a schedule coefficient that varies with , and is the total sampling timesteps. A noise estimation network plays a critical role in approximating the noise distribution in the diffusion process. Specifically, it aims to minimize the discrepancy between the predicted noise term and the actual noise . In most current works, the noise estimation network adopts the DiT architecture, where the predicted noise function can be further reformulated as:
@@
Classifier-Free Guidance (CFG) has proven to be a powerful technique for improving the fidelity of generated videos in diffusion models. During the sampling process, CFG generates two distinct predictions: the conditional output conditioned on the input context , and the unconditional output derived from the empty/negative prompt . The final denoised output is:
@@ 3.2. Analysis and Motivation
Potential of three-level cache integration. We implement different granularity caching methods at each diffusion timestep, and adopt the LPIPS (Zhang et al., 2018) metric to calculate the end-to-end similarity with the original video. The smaller the LPIPS value, the higher the similarity. As shown in Figure 3, the optimal cache granularity varies dynamically across timesteps, suggesting that adaptive cache selection could enhance generation quality.
@@
3.3. Context-aware Cache Triggering
As the early diffusion stage is responsible for overall framework sketching, it is highly sensitive to interference at this time, as described in SRDiffusion (Cheng et al., 2025). This phenomenon can also be validated in Figure 2, where in the initial diffusion process, the three-level redundancy is relatively low. Therefore, we perform full computation in the initial diffusion stage, and calculate between the output of the current timestep and that of the previous timestep , and we call this stage as the warm up stage. When is smaller than the predefined threshold , established in the offline profiling process (detailed later), the warm up phase ends and enters the cache enabled phase.
@@ 3.3. Context-aware Cache Triggering
3.4. Adaptive Hybrid Cache Decision
@@
After identifying cache enabled timesteps, it is crucial to determine the specific caching granularity for a certain timestep within the three-level caching. The limitation of prior works (Liu et al., 2024, 2025; Ma et al., 2025) is their exclusive focus on similarity of different cache methods while neglecting their differential impact on accuracy. To assess the accuracy impact of different caching methods, we generate Gaussian distribution with predefined statistical parameters (mean , standard deviation ) and calculate the distance between perturbed and original outputs. It should be noted that and should align with the real parameters derived from the three-level caching. Therefore, we firstly profile prompts and examine the and of the distance tensor between the real output and the cached output. As shown in Figure 6, the value of the distance tensor is inconsistent across different prompts, and also varies at different timesteps. However, its values are fixed within a certain range. Therefore, we set the value as the average of value across all timesteps. In addition, for the value, except for the initial timesteps, it remains stable at a fix value. Therefore, we set the value as the average of the value of the last 40 timesteps. We exclude the first 10 timesteps from analysis since these initial timesteps correspond to the warm up phase without applying caching. Consequently, their outputs lack statistical representativeness for evaluating the proposed methodology.
@@
After determining the and value, we examine the relative L1 distance between the perturbed output and the original output. We examine for all timesteps and plot the results in Figure 5(a). The smaller the distance, the smaller the impact of the interference on the actual results. This value can indicate the accuracy impact of performing a specific level cache at a specific timestep. From Figure 5(a), it can be seen that the impact value of step level and cfg level remain relatively stable after the warm up phase. However, the block level value exhibits time-dependent characteristics, where its initial value is low, and then increases in the later diffusion stage. In addition, the compare across three levels reveals that, the cfg level interference exerts a substantially greater influence, with its value exceeding both step level and block level interference by an order of magnitude. We present the visualized results and similarity metrics in Figure 5(b). It can be seen that the video quality generated using cfg level interference is poor, indicating that this interference has a significant impact on the final results, which is consistent with the conclusion in Figure 5(a). Therefore, the values presented in Figure 5(a) are employed as quantitative evaluation metrics for assessing accuracy impact, with the impact of step level and cfg level remaining constant (, ), while the block level impact exhibit as a time-dependent and index-dependent function (). Thus, the offline Gaussian perturbation estimates the quantified impact value for each cache type using and fit from measured cache-induced differences.
@@
where i is one of the pre-defined block indices. To prevent getting stuck in the same caching granularity within a cache interval, we introduce a penalty strategy: if a caching granularity is used in one timestep, it is disabled in the next timestep. This proactive approach ensures flexibility and robustness, enabling systems to avoid local optima.
3.5. Unify it together: MixCache Framework
@@
According to Figure 2, the redundancy level of different models varies. Given a specific model, MixCache first performs offline profiling to determine key hyper-parameters. We run 100 prompts through the diffusion process, recording at each timestep. Once stabilizes, which is identified when its value remains below a certain value compared to the previous five steps, we compute the average as . The maximum and minimum deviations from after stabilization are set as and . These values are averaged across all prompts and rounded as needed. The profiling also determines and for Gaussian interference, used to assess impact at different granularities. This process is performed once per model and enables efficient runtime deployment without extra overhead.
After the offline profiling process, it enters the runtime inference phase, and the execution flow is formalized in Algorithm 1. During runtime, one generation is partitioned into two distinct phases: the warm up phase and the cache enabled phase. In the warm up phase, the system exclusively executes full computations to maintain video quality, and this phase is ended by the context-aware cache triggering strategy to determine which timestep to enable caching (line 5). Once into the cache enabled phase, MixCache uses the adaptive hybrid cache decision strategy to determine which of the three-level caching methods to use. There are two core mechanisms are activated: (1) the scaling strategy to regulate the frequency of cache interval (line 11), and (2) the adaptive hybrid cache decision strategy (line 23-24) that dynamically determines the optimal caching granularity for each timestep.
4. Experiments
@@
4.1. Settings
Base Models and Baselines. We evaluate MixCache on Wan2.1 14B (Wang et al., 2025), HunyuanVideo (Kong et al., 2024) and CogVideoX 5B (Yang et al., 2024b). For baseline methods, we choose Teacache (Liu et al., 2024), FasterCache (Lv et al., 2024), BlockDance (Zhang et al., 2025) and PAB (Zhao et al., 2025), all of which are specifically designed to accelerate DiT models through caching. Among them, Teacache adopts step level cache, FasterCache combines cfg level and block level cache, and BlockDance and PAB employ block level cache. We adjust the parameters of each baseline to achieve a balance between speed and quality. Specifically, Teacache0.1 indicates that the value of L1_distance_thresh from the open-source implementation is set to 0.1, and the implementation of PAB is adopted on HuggingFace Diffusers, where PAB denotes a configuration of block_skip_range=8 and timestep_skip_range=. For BlockDanceN, following the official setting, we set the cache block index as 20 and set the first 40% of denoising steps as warm up steps that disable caching, and evenly divide the remaining 60% of denoising steps into several groups, each comprising steps.
Evaluation Metrics. To evaluate the quality of video generation, we employ VBench (Huang et al., 2024), a widely-adopted comprehensive benchmarking suite for evaluating video generation. Based on VBench standard prompt set, we use Qwen2.5-14B-Instruct (Yang et al., 2024a) to extend all prompts to enhance the video quality, and generate 5 videos with different seeds for each prompt. In addition, we report LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004) and PSNR for quality comparison. For efficiency evaluation, we quantify the average inference latency per prompt as the primary performance metric.
Experiment Settings. For the main results, we generated 4720 videos for each set of results based on VBench. These were processed in data parallel across a public cloud instance with 64 NVIDIA A800 (80GB) GPUs, and the whole experiment takes nearly a month to complete. Except the main results that use all prompts of Vbench, we randomly sample 200 prompts from VBench to conduct ablation experiments, each using 1 seed for generation.
Implementation Details. All experiments conduct full attention inference. We pre-defined block index as (10,20,30) as block caching candidates, which performs well for most models. The profiled hyper-parameters and (, ) for Wan and HunyuanVideo are 0.1, (0.05, 0.1), respectively, and those for CogVideoX are 0.1, (0.3, 0.4) respectively.
4.2. Main Results
Quantitative Comparison. Table 1 presents a quantitative comparison of MixCache with baselines in terms of efficiency and visual quality. The results highlight that, MixCache consistently demonstrates robust acceleration efficiency and maintains visually compelling quality across diverse base models, baselines, and resolutions.
Visual Comparison. Figure 7 compares the videos generated by MixCache against those by the original model and baselines. The results demonstrate that MixCache can effectively preserve the original semantics and fine details.
4.3. Ablation Study
In the ablation study, we investigate the effects of scaling strategy, penalty strategy and hybrid three-level caching respectively in terms of the quality and efficiency. The quantitative results are shown in Table 2 and the corresponding visualization results are presented in Figure 8. Among them, “hybrid” refers to using the adaptive hybrid cache decision strategy to determine the cache method for each cache enabled step, while “step/cfg/block only” refers to using only a specific caching method in all cache enabled steps. “N=4” represents a fixed cache interval of 4, and “” represents enabling the scaling strategy that prioritizes efficiency. It can be seen that MixCache outperforms all ablation experiments in both quality and efficiency metrics, indicating that combining dynamic scaling, penalty strategy and hybrid three-level caching can achieve better generation quality and inference efficiency.